# The EU AI Act, Stakeholder Needs, and Explainable AI: Aligning Regulatory Compliance in a Clinical Decision Support System

### v2

@@ 6.1 Roles, Responsibilities and Needs
As laid out in Section˜2.2, the XAI community is concerned with stakeholder desiderata (langer_what_2021) while the AI Act concerns stakeholder responsibilities (see Section˜2.1). While there is an overlap in the main stakeholder roles introduced by XAI and the AI Act, there are also differences. Where the XAI community is satisfied by the role of regulator (langer_what_2021), the AI Act has multiple roles. There is the national authority assigned to conduct market surveillance of AI systems, the notified body that inspects to what extent an AI system and the developer’s organization conform to the requirements and the EU Commission who all have a say in how the AI Act is to be implemented and enforced. Even a court could serve as a regulator if the system developer and the national authority have different opinions. In the case of our CDSS, the system developer can even take the XAI role of regulator. This can seem counterintuitive, but it is one way of interpreting the conformity procedure based on self-assessment.
On the other side, the XAI community pays explicit attention to the end user and the affected parties (langer_what_2021), while the AI Act trusts the provider and the deployer to take care of those perspectives. As such, the stakeholder roles can be seen as lacking a common ground, or as complementing each other. That leaves ample room for the XAI community to define their stakeholder roles in relation to the responsibilities enforced by the AI Act.
Another observation is that while langer_what_2021 treat individuals “as representative members of specified stakeholder classes”, our analysis mandates us to distinguish between affected parties as concrete individuals and as an abstract collective due to the responsibilities of the deployer. Art. 26 in the AI Act defines requirements on the usage of the AI system on an individual level at run-time, while Art. 27 poses requirements on performing risk assessments of the AI system in relation to its intended usage pre-launch in respect to the affected parties in general. This is visualized in Figure˜2(a) as two different responsibilities (arrows) between the deployer and the affected parties.
Key Insight (KI1) The XAI community should define their roles in relation to the responsibilities enforced by the AI Act.
@@ 6.2 Bridging the AI Act’s Gap
The AI Act deliberately excludes individual patients’ will, focusing instead on roles relevant to product safety. Interestingly, the XAI perspective does not give the patient any more agency than the AI Act does. Thus, both the AI Act and XAI have roles that act on behalf of the affected party, which by itself lacks the ability, mandate, or obligation to change the functionality of the AI system. However, XAI emphasizes the needs of affected parties in relation to the user, potentially bridging the gap between end users and affected parties, which the AI Act perspective lacks (highlighted as blue arrow in Figure˜2(a)).
Notably, the AI Act does not explicitly define stakeholder roles, instead referring to “relevant stakeholders”, “external stakeholders”, “multi-stakeholder” etc. It is therefore our analysis that defines Deployer, Competent Authority, and Provider as stakeholder roles. Regarding the equivalence to the XAI roles of User and Affected parties, the AI Act refers to the former as “staff” (the collective that shall be informed, trained and have responsibilities regarding AI systems according to Art. 17 on quality management systems), “affected workers” (those who will use the CE-marked system in relation to Art. 26 and the impact on individuals’ rights) and “natural persons” (those who have the adequate training for using a CE-marked AI system, Art. 26). The latter role is referred to as “categories of natural persons or groups of persons” (impacted by CE-marked systems under Art. 27). As the AI Act lacks systematic terminology for stakeholder roles, we use Natural person and Natural Person(s) to emphasize human involvement (natural person is legal terminology for human) rather than legal persons (legal terminology for organizations, companies, authorities, etc.).
Key Insight (KI2) XAI enhances the AI Act by focusing on the relation between the affected parties and the user and their needs.
@@ 6.3 From Regulator to Affected Parties—and back
The AI Act mandates respecting the fundamental rights of affected parties, but does not specify technical methods or thresholds. This opens a research trajectory for the XAI community on how to achieve appropriate methods and reasonable cut-off points. As an example, how certain does a recommendation system need to be about key-indicators, such as the LOS, that support the decision which patient is moved from the ward to make room for another patient? How do we calculate that certainty, and how is it represented to physicians, hospital management, regulators, and affected parties? For this to be relevant in a legal context, the future contributions of the XAI community need to be understood and accepted by the diverse group of actors that can take the stakeholder roles of deployer, developer and regulator (see Figure˜2(b)). This aligns with pavlidis_unlocking_2024, who emphasizes that XAI must operate in compliance settings to meet regulatory requirements. In our case, it led us to target interpretability, usability and trust as key needs to address.
The AI Act introduces operational activities like a data lifecycle and risk management aspects (e.g., the delegation of responsibilities), without specifying methods. Here, the XAI community has an opportunity to delimit what is within its interests to contribute towards, but also new aspects of XAI development and usage for future contributions.
In our case, FRIA must be conducted before system deployment, requiring the deployer to assess the impact and risks to patients’ fundamental rights (Art. 27, which relates to Fairness). This links back to the provider’s obligations. The FRIA uses information from the AI-system provider (Art. 13) and the provider’s risk assessment (Art. 9). Technical documentation and logs help the provider conduct the FRIA and assist the user in evaluating output relevance, impacting XAI desiderata like trust and acceptance.
Key Insight (KI3) XAI facilitates compliance with the AI Act by offering essential documentation for FRIA, thereby enhancing trust and acceptance among end-users.
@@ 7 Discussion and future research
From the XAI perspective, the starting point of our analysis is the obligations of the deployer. The persons assigned the oversight or usage of the system need to have the right competence, authority, and support so that they can assess if the output is reasonable given the input and current state of the ICU (Art. 26 of the AI Act, which relates to the XAI term of Interpretability). As described in Section˜3.2, the CDSS provides key indicator recommendations to support the clinical decision, which patients can be moved to another ward. While XAI can provide measures for explaining how the recommendation came to be, it does not assess whether it is right or wrong. Operating the ward still requires human judgment to assess which recommendations to follow and when alternative routes are more appropriate (see, e.g., (schoeffer_explanations_2024; gomez_human-ai_2025) on Human-AI Collaboration).
@@
Yet, there remains much to do. Since the XAI method choices reflect what happens between the physician and the patient (novice user and affected party), they need to be percolated upwards to the hospital management (deployer) of the AI system and captured as requirements, consequently fulfilled by the system developer (developer) of the AI system. Finally, the design choices and their implementation need to be understood and accepted by the regulator (see Figure˜2(b)). XAI can serve as an instrument to build trust at all levels in relation not only to AI as technology but to a specific AI system – the regulator trusts the AI system to be fit for purpose, the deployer and provider trust the AI system to be fit for purpose and the user and the affected parties trust the AI system to be fit for purpose. A similar reasoning is applicable for the usability of the AI system — it needs to be usable for both the end user and the deployer. We foresee a number of studies to further explore the relationship between the levels but also between XAI as a tool and the legal requirements of the AI Act and how they are further detailed in the upcoming harmonized standards.
Here, it is worth pointing to an emerging discussion within the academic community on the risks of introducing AI in the healthcare system, specifically in the relationship between clinician and patient and the liability concerns in relation to perceived mistreatment (see, e.g., (solaiman_regulating_2024; Duffourc2023; Cartolovni2023)). XAI could both serve as a means to lessen the risk of litigation, but also, in itself, poses a risk if not implemented properly. This is exemplified by weller_transparency_2022, who states that explanations can lead to unsubstantiated trust to the lay user, and hence have a negative impact on affected parties. If XAI is to realize its potential in relation to the AI Act it is important that it is not applied to “XAI-wash” unsuitable systems (see e.g., rudin_stop_2019 who criticizes the use of XAI methods that may not be suitable in high-stakes decisions which are nonetheless justified using XAI), but for a genuine wish to make AI a suitable technology for human needs.
@@ 8 Conclusion
As the AI Act enters into force, AI-based systems are now regulated under comprehensive product safety rules. Therefore, the AI Act refers to stakeholder obligations without putting the end-users and their needs at the center of attention. On the other hand, the research around XAI focuses on stakeholder desiderata. By analyzing a CDSS under development in relation to the AI Act and by analyzing the stakeholders from different perspectives, we map stakeholders’ desiderata to the requirements of the AI Act. In doing so, we discuss the role of XAI in the AI Act and make the stakeholder role of the (novice) user and the subsequent relation to the affected parties explicit.
Especially in high-risk areas like healthcare, we emphasize the critical role of human oversight and competence in deploying AI systems to ensure trustworthy outputs and appropriate decisions. We highlight the necessity of conducting FRIA and engaging stakeholders in selecting tailored XAI methods to build trust and acceptance. We see XAI as a chance to initiate a dialogue on the effective implementation of different elements of the AI Act and as an instrument to build trust in a specific AI system at all stakeholder levels.
Statements and Declarations
Competing Interests
The authors have no relevant financial or non-financial interests to disclose.

### v3

@@ 6.1 Roles, Responsibilities and Needs
While there is an overlap in the main stakeholder roles introduced by Langer et al. (2021) and the AI Act, there are also differences. Where Langer et al. (2021) taxonomy includes a single regulator role, the AI Act has multiple roles. There is the national authority assigned to conduct market surveillance of AI systems, the notified body that inspects to what extent an AI system, the developer’s organization conforming to the requirements, and the EU Commission who have a say in how the AI Act is to be implemented and enforced. Even a court could serve as a regulator if the system developer and the national authority have different opinions. In the case of our CDSS, the system developer can even take the XAI role of regulator. This can seem counterintuitive, but it is one way of interpreting the conformity procedure based on self-assessment.
On the other hand, Langer et al. (2021) pay explicit attention to the end user and the affected parties, while the AI Act trusts the provider and the deployer to take care of those perspectives. As such, the stakeholder roles can be seen as lacking a common ground, or as complementing each other. That leaves ample room for the XAI community to define their stakeholder roles in relation to the responsibilities enforced by the AI Act.
Another observation is that while Langer et al. (2021) treat individuals “as representative members of specified stakeholder classes”, our analysis mandates us to distinguish between affected parties as concrete individuals and as an abstract collective due to the responsibilities of the deployer. Art. 26 in the AI Act defines requirements on the usage of the AI system on an individual level at run-time, while Art. 27 poses requirements on performing risk assessments of the AI system in relation to its intended usage pre-launch in respect to the affected parties in general. This is visualized in Figure˜2(a) as two different responsibilities (arrows) between the deployer and the affected parties.
Key Insight (KI1) XAI practitioners must define their roles across AI Act responsibilities — moving from research observers to active participants in regulatory implementation.
@@ 6.2 Bridging the AI Act’s Gap
The AI Act deliberately excludes the will of the affected parties (in our case, individual patients), focusing instead on roles relevant to product safety. Both the AI Act and XAI have roles that act on behalf of the affected party. In the AI act, the affected party lacks the ability, mandate, or obligation to change the functionality of the AI system. In contrast, XAI emphasizes the needs of affected parties in relation to the user, potentially bridging the gap between end users and affected parties, which the AI Act perspective lacks (highlighted as blue arrow in Figure˜2(a)).
Notably, the AI Act does not explicitly define stakeholder roles, instead referring to “relevant stakeholders”, “external stakeholders”, “multi-stakeholder” etc. It is therefore our analysis that defines Deployer, Competent Authority, and Provider as stakeholder roles. Regarding the equivalence to the XAI roles of User and Affected parties, the AI Act refers to the former as “staff” (the collective that shall be informed, trained and have responsibilities regarding AI systems according to Art. 17 on quality management systems), “affected workers” (those who will use the CE-marked system in relation to Art. 26 and the impact on individuals’ rights) and “natural persons” (those who have the adequate training for using a CE-marked AI system, Art. 26). The latter role is referred to as “categories of natural persons or groups of persons” (impacted by CE-marked systems under Art. 27). As the AI Act lacks systematic terminology for stakeholder roles, we use natural person(s) to emphasize human involvement (natural person is legal terminology for human) rather than legal persons (legal terminology for organizations, companies, authorities, etc.).
In contrast, stakeholder-oriented XAI methods explicitly consider the affected parties, for instance by personalizing explanations and thus directly addressing the perspectives of individuals affected by AI-based systems (Uth et al., 2026; Quah et al., 2025). By customizing explanations based on their characteristics (Schneider and Handali, ), they become understandable across different levels of domain and AI expertise. This shared understanding of AI explanations facilitates dialogue among and between different user groups and affected parties (e.g., clinicians and patients).
Key Insight (KI2) XAI fills a critical gap in the AI Act: while regulations address provider-deployer obligations, XAI makes the user-to-affected-party relationship explicit and actionable.
@@ 6.3 From Regulator to Affected Parties — and back
The AI Act mandates respecting the fundamental rights of affected parties. In our case, a FRIA must be conducted before system deployment, requiring the deployer to assess the impact and risks to patients’ fundamental rights (Art. 27, which relates to Fairness). This links back to the provider’s obligations. The FRIA uses information from the AI-system provider (Art. 13) and the provider’s risk assessment (Art. 9). Technical documentation and logs help the provider conduct the FRIA and assist the user in evaluating output relevance, impacting XAI desiderata like trust and acceptance. However, the AI Act does not specify technical methods or thresholds.
This opens a research trajectory for the XAI community on how to achieve appropriate methods and reasonable cut-off points. As an example, how certain does a recommendation system need to be about key-indicators, such as the LOS, that support the decision which patient is moved from the ward to make room for another patient? How do we calculate that certainty, and how is it represented to physicians, hospital management, regulators, and affected parties? For this to be relevant in a legal context, the future contributions of the XAI community need to be understood and accepted by the diverse group of actors that can take the stakeholder roles of deployer, developer and regulator (see Figure˜2(b)). This aligns with Pavlidis (2024), who emphasizes that XAI must operate in compliance settings to meet regulatory requirements. In our case, it led us to target interpretability, usability and trust as key needs to address.
The AI Act introduces additional operational activities like a data lifecycle and risk management aspects (e.g., the delegation of responsibilities), without specifying methods. Here, the XAI community has an opportunity to delimit what is within its interests to contribute towards, but also new aspects of XAI development and usage for future contributions.
Key Insight (KI3) XAI simultaneously facilitates compliance with the AI Act by offering essential documentation for FRIA and builds multi-level trust — from regulators to affected parties.
@@ 7 Discussion and future research
From the XAI perspective, the starting point of our analysis is the obligations of the deployer. The persons assigned the oversight or usage of the system need to have the right competence, authority, and support so that they can assess if the output is reasonable given the input and current state of the ICU (Art. 26 of the AI Act, which relates to the XAI term of Interpretability). As described in Section˜3.2, the CDSS provides key indicator recommendations to support the clinical decision which patients can be moved to another ward. While XAI can provide measures for explaining how the recommendation came to be, it does not assess whether it is right or wrong. Operating the ward still requires human judgment to assess which recommendations to follow and when alternative routes are more appropriate (see, e.g., (Schoeffer et al., 2024; Gomez et al., 2025) on Human-AI Collaboration).
@@
Yet, there remains much to do. Since the XAI method choices reflect what happens between the physician and the patient (novice user and affected party), they need to be percolated upwards to the hospital management (deployer) of the AI system and captured as requirements, consequently fulfilled by the system developer (developer) of the AI system. Finally, the design choices and their implementation need to be understood and accepted by the regulator (see Figure˜2(b)). XAI can serve as an instrument to build trust at all levels in relation not only to AI as technology but to a specific AI system — the regulator trusts the AI system to be fit for purpose, the deployer and provider trust the AI system to be fit for purpose and the user and the affected parties trust the AI system to be fit for purpose. A similar reasoning is applicable for the usability of the AI system — it needs to be usable for both the end user and the deployer. We foresee a number of studies to further explore the relationship between the levels but also between XAI as a tool and the legal requirements of the AI Act and how they are further detailed in the upcoming harmonized standards.
Here, it is worth pointing to an emerging discussion within the academic community on the risks of introducing AI in the healthcare system, specifically in the relationship between clinician and patient and the liability concerns in relation to perceived mistreatment (see, e.g., (Solaiman and Malik, 2024; Duffourc2023; Cartolovni2023)). XAI could both serve as a means to lessen the risk of litigation, but also, in itself, poses a risk if not implemented properly. This is exemplified by Weller (2022), who states that explanations can lead to unsubstantiated trust to the lay user, and hence have a negative impact on affected parties. If XAI is to realize its potential in relation to the AI Act it is important that it is not applied to “XAI-wash” unsuitable systems (see e.g., (Rudin, 2019) who criticizes the use of XAI methods that may not be suitable in high-stakes decisions which are nonetheless justified using XAI), but for a genuine wish to make AI a suitable technology for human needs.
@@ 8 Conclusion
As the AI Act enters into force, AI-based systems are now regulated under comprehensive product safety rules. The AI Act mandates organizational obligations while underrepresenting end-users — creating a critical opportunity for human-centered XAI. By analyzing a CDSS under development in relation to the AI Act and by analyzing the stakeholders from different perspectives, we map stakeholders’ desiderata to regulatory requirements, making the role of the (novice) user and the relation to affected parties explicit. We demonstrate how human-centered XAI can bridge the gap between legal compliance and stakeholder needs.
Our analysis reveals three key opportunities for the XAI community. First, human-centered XAI practitioners must actively define their roles across AI Act responsibilities — moving from research observers to active participants in regulatory implementation. Second, human-centered XAI fills a critical gap: while regulations address provider-deployer obligations, XAI makes the user-to-affected-party relationship explicit and actionable. Third, stakeholder-tailored XAI simultaneously enables regulatory compliance (FRIA, documentation, risk management) and builds multi-level trust — from regulators to affected parties.
We call on the XAI community to seize this moment: while regulatory compliance requires multiple components — from harmonized standards to risk management frameworks — human-centered XAI plays a unique role in ensuring transparency mechanisms address genuine stakeholder needs. This bidirectional relationship where regulations reshape XAI research priorities while XAI practitioners shape transparency implementation offers an opportunity to ensure that those affected by AI systems and their regulation have their interests represented in regulatory practice.